OpenStudy (anonymous):

If the postal charges for a package are 62 cents for the first five ounces and 8cents for each additional ounce, what is the weight of a package for which thecharges are $1.66? (Assume there are 16 ounces in one pound)

OpenStudy (saifoo.khan):

im not sure/

OpenStudy (saifoo.khan):

Let me retry.

OpenStudy (ajprincess):

For the first 5 ounces the charge is 62 cents subtracting it from the total charge, we get 1.66-0.62 =1.04 the charge for additional ounce=8cents To see hw many additional ounces are added divide 1.04 by 0.08, then we get 1.04/0.08=13 So the total weight of the parcel=13+5 =18ounces 16 ounces=1 pound The weight=1pound 2ounces.